During 2019, Jordan Company was subject to the Alaska state unemployment tax of 4.2%. The company's taxable wages for FUTA were $86,700 and for SUTA, $171,000. Compute the following; round your answers to the nearest cent.

1.SUTA tax that Jordan Company would pay to the state of Alaska
$

2.Net FUTA tax for 2019
$

3.Amount of employees' unemployment tax for 2019

$
Answer #1

1.SUTA tax that Jordan Company would pay to the state of Alaska $

$171,000*4.2/100 = $ 7182

2.Net FUTA tax for 2019 $

Net FUTA Tax rate = 6-4.2 = 1.8%

Tax Payable = 86700*1.8/100 = $ 1561

3.Amount of employees' unemployment tax for 2019

= 7182+1561 = $ 8743